# Explainable Multi-modal Time Series Prediction with LLM-in-the-Loop

### v1

@@ 3.3.2 Iterative Context Refinement via Reflective Feedback
Algorithm 1 TimeXL: Explainable Multi-modal Time Series Prediction with LLM Agents
Inputs: Multi-modal time series $(\boldsymbol{x},\boldsymbol{s},\boldsymbol{y})$, prototype-based encoder $\mathcal{M}_{\mathrm{enc}}$, prediction agent $\mathcal{M}_{\mathrm{pred}}$, reflection agent $\mathcal{M}_{\mathrm{refl}}$, refinement agent $\mathcal{M}_{\mathrm{refine}}$, fusion parameter $\alpha$, max iteration $\tau$, improvement evaluation $\mathrm{Eval(\cdot)}$ based on metrics
Training:
Initialize $\boldsymbol{s}_{0}=\boldsymbol{s}$, $i=0$, $\boldsymbol{\hat{y}_{\mathrm{all}}}=\{\}$
while $\mathrm{Eval}(\boldsymbol{\hat{y}}_{\mathrm{all}},\boldsymbol{y})$ not pass or iteration $i<\tau$ do
       Train $\mathcal{M}_{\mathrm{enc}}$ using multi-modal data $\mathcal{D}_{i}=\{(\boldsymbol{x},\boldsymbol{s}_{i},\boldsymbol{y}),\cdots\}$
Infer explainable prediction $\boldsymbol{\hat{y}}_{\mathrm{enc}},\mathbf{expl}_{\boldsymbol{s}_{i}}=\mathcal{M}_{\mathrm{enc}}(\boldsymbol{x},\boldsymbol{s}_{i})$
Infer LLM prediction $\boldsymbol{\hat{y}}_{\mathrm{LLM}}=\mathcal{M}_{\mathrm{pred}}(\boldsymbol{s}_{i},\mathbf{expl}_{\boldsymbol{s}_{i}})$
Fuse prediction $\boldsymbol{\hat{y}}=\alpha\boldsymbol{\hat{y}}_{\mathrm{enc}}+(1-\alpha)\boldsymbol{\hat{y}}_{\mathrm{LLM}}$
Generate reflection $\mathrm{Refl}=\mathcal{M}_{\mathrm{refl}}(\boldsymbol{y},\boldsymbol{\hat{y}}_{\mathrm{LLM}},\boldsymbol{s}_{i})$
Refine text based on reflection $\boldsymbol{s_{i+1}}=\mathcal{M}_{\mathrm{refine}}(\mathrm{Refl},\boldsymbol{s}_{i})$
Append $\boldsymbol{\hat{y}}$ to $\boldsymbol{\hat{y}}_{\mathrm{all}}$
@@
return $\mathcal{M}_{\mathrm{enc}}$, $\mathrm{Refl}$, $\boldsymbol{s}_{i+1}$ Validation and Testing:
Refinement based on reflection $\boldsymbol{s^{\prime}}=\mathcal{M}_{\mathrm{refine}}(\mathrm{Refl},\boldsymbol{s})$
Infer explainable prediction $\boldsymbol{\hat{y}}_{\mathrm{enc}},\mathbf{expl}_{\boldsymbol{s^{\prime}}}=\mathcal{M}_{\mathrm{enc}}(\boldsymbol{x},\boldsymbol{s^{\prime}})$
Infer LLM prediction $\boldsymbol{\hat{y}}_{\mathrm{LLM}}=\mathcal{M}_{\mathrm{pred}}(\boldsymbol{s^{\prime}},\mathbf{expl}_{\boldsymbol{s^{\prime}}})$
Fuse prediction $\boldsymbol{\hat{y}}=\alpha\boldsymbol{\hat{y}}_{\mathrm{enc}}+(1-\alpha)\boldsymbol{\hat{y}}_{\mathrm{LLM}}$
Given the prediction $\boldsymbol{\hat{y}}_{\mathrm{LLM}}$ generated by the prediction agent $\mathcal{M}_{\mathrm{pred}}$, the reflection agent $\mathcal{M}_{\mathrm{refl}}$ aims to understand the reasoning behind the implicit prediction logic of $\mathcal{M}_{\mathrm{pred}}$. Specifically, it generates a reflective feedback, $\mathrm{Refl}$, by analyzing the input text $\boldsymbol{s}$ and its prediction $\boldsymbol{\hat{y}}_{\mathrm{LLM}}$, against the ground truth $\boldsymbol{y}$, to provide actionable insights for refinement, i.e., $\mathrm{Refl}=\mathcal{M}_{\mathrm{refl}}(\boldsymbol{y},\boldsymbol{\hat{y}}_{\mathrm{LLM}},\boldsymbol{s})$. Guided by the feedback, the refinement agent $\mathcal{M}_{\mathrm{refine}}$ refines the previous text $\boldsymbol{s}_{i}$ into $\boldsymbol{s}_{i+1}$ by selecting and emphasizing the most relevant content, ensuring that important patterns are appropriately contextualized, which is similar to how a domain expert would perform, i.e., $\boldsymbol{s_{i+1}}=\mathcal{M}_{\mathrm{refine}}(\mathrm{Refl},\boldsymbol{s}_{i})$. The prompts for querying $\mathcal{M}_{\mathrm{refl}}$ and $\mathcal{M}_{\mathrm{refine}}$ are provided in Figures 1415 1617, and discussed in Appendix D.
We finally integrate the refinement via reflection into the optimization loop of our proposed TimeXL, which is summarized in Algorithm 1. Once the textual context is improved, it is used to retrain the multi-modal prototype-based encoder $\mathcal{M}_{\mathrm{enc}}$ for the next iteration. As such, the explanation (e.g., quality of the prototypes) and predictive performance of $\mathcal{M}_{\mathrm{enc}}$ can be improved through this iterative process. Consequently, the prediction agent $\mathcal{M}_{\mathrm{pred}}$ could yield better prediction with more informative inputs, further enhancing the accuracy of $\boldsymbol{\hat{y}}$. We evaluate the trajectory of predictive performance and terminate the iteration if at least an improvement is observed ($\mathrm{Eval(\cdot)}$ pass) when max iteration is reached. Note that, in the testing phase, we use the reflection $\mathrm{Refl}$ generated in the best training iteration (evaluated on validation set) to guide $\mathcal{M}_{\mathrm{refine}}$ for context refinement, mimicking how an optimized deep model is applied to testing data.
@@ 4 Experiments
Table 1: The F1 score (F1) and AUROC (AUC) for TimeXL and state-of-the-art baselines on real-world multi-modal time series datasets.
@@
DLinear (Zeng et al., 2023) 0.540 0.660 0.255 0.485 0.393 0.500 0.419 0.388
Autoformer (Wu et al., 2021) 0.546 0.590 0.565 0.747 0.774 0.918 0.683 0.825
Crossformer (Zhang & Yan, 2023) 0.500 0.594 0.571 0.775 0.924 0.984 0.737 0.913
TimesNet (Wu et al., 2023) 0.494 0.594 0.538 0.756 0.794 0.867 0.765 0.944
iTransformer (Liu et al., 2024c) 0.541 0.650 0.600 0.783 0.861 0.931 0.791 0.963
TSMixer (Chen et al., 2023a) 0.488 0.534 0.465 0.689 0.770 0.797 0.808 0.931
FreTS (Yi et al., 2024) 0.623 0.688 0.546 0.737 0.887 0.950 0.751 0.762
PatchTST (Nie et al., 2023) 0.592 0.675 0.604 0.795 0.841 0.934 0.695 0.928
LLMTime (Gruver et al., 2023) 0.587 0.657 0.315 0.498 0.802 0.817 0.769 0.803
PromptCast (Xue & Salim, 2023) 0.499 0.365 0.418 0.607 0.727 0.768 0.696 0.871
OFA (Zhou et al., 2023) 0.501 0.606 0.512 0.745 0.774 0.879 0.851 0.977
Time-LLM (Jin et al., 2024) 0.613 0.699 0.589 0.792 0.671 0.864 0.733 0.912
TimeCMA (Liu et al., 2025) 0.636 0.731 0.559 0.727 0.729 0.828 0.693 0.843
MM-iTransformer (Liu et al., 2024a) 0.608 0.689 0.605 0.793 0.926 0.986 0.901 0.990
MM-PatchTST (Liu et al., 2024a) 0.621 0.718 0.619 0.812 0.863 0.968 0.780 0.929
TimeCAP (Lee et al., 2025) 0.668 0.742 0.611 0.801 0.954 0.983 0.942 0.988
@@ 4.1 Experimental Setup
Datasets. We evaluate methods on four multi-modal time series datasets from three different real-world domains, including weather, finance, and healthcare. The detailed data statistics are summarized in Table 3 of Appendix A.1. The weather dataset contains meteorological reports and the hourly time series records of temperature, humidity, air pressure, wind speed, and wind direction in New York City. The task is to predict if it will rain in the next 24 hours, given the last 24 hours of weather records and summary. The finance dataset contains the daily record of the raw material prices together with 14 related indices from January 2017 to July 2024. Given the last 5 business days of stock price data and news, the task is to predict if the target price will exhibit an increasing, decreasing, or neutral trend on the next business day. The healthcare datasets contain Test-Positive (TP) and Mortality (MT). The Test-Positive dataset consists of the weekly records and healthcare reports of the number of positive specimens for Influenza A and B. The task is to predict if the percentage of respiratory specimens testing positive in the upcoming week for influenza will exceed the average value, given the records and summary in the last 20 weeks. Similarly, the Mortality dataset contains the weekly records and reports of influenza and pneumonia deaths. The task is to predict if the mortality ratio from influenza and pneumonia will exceed the average value, given the records and summary in the last 20 weeks.
Baselines, Evaluation Metrics and Setup We compare TimeXL with state-of-the-art baseline methods for time series prediction. These baselines includes Autoformer (Wu et al., 2021), Dlinear (Zeng et al., 2023), Crossformer (Zhang & Yan, 2023), TimesNet (Wu et al., 2023), PatchTST (Nie et al., 2023), iTransformer (Liu et al., 2024c), FreTS (Yi et al., 2024), TSMixer (Chen et al., 2023a) and LLM-based methods like LLMTime (Gruver et al., 2023), PromptCast (Xue & Salim, 2023), OFA (Zhou et al., 2023), Time-LLM (Jin et al., 2024) and TimeCMA (Liu et al., 2025), where LLMTime and PromptCast don’t need fine-tuning. While these methods are primarily used for time series prediction with continuous values, they can be easily adapted for discrete value prediction. We also evaluate the multi-modal time series methods. Besides the Time-LLM and TimeCMA where input text is used for embedding reprogramming and alignment, we also evaluate Multi-modal PatchTST and Multi-modal iTransformer from (Liu et al., 2024a), as well as TimeCAP (Lee et al., 2025). We evaluate the discrete prediction via F1 score and AUROC (AUC) score, due to label imbalance in real-world time series datasets. We split all datasets for training/validation/testing by a ratio of 6/2/2. We alternate different embedding methods for texts based on its average length, where we use Bert (Kenton & Toutanova, 2019) as the embedding model for weather and healthcare datasets, and sentence transformer (Reimers & Gurevych, 2019) for finance dataset.
@@ 4.2 Performance Evaluation
The results of predictive performance are shown in Table 1. It is notable that multi-modal methods generally outperform time series methods across all datasets. These methods include LLM methods (e.g., Time-LLM, TimeCMA) that leverage text embeddings to enhance time series predictions. Moreover, the multi-modal variants (MM-iTransformer and MM-PatchTST) improve the performance of state-of-the-art time series methods, suggesting the benefits of integrating real-world contextual information. Besides, TimeCAP integrates the predictions from both modalities, further improving the predictive performance. TimeXL constantly achieves the highest F1 and AUC scores, consistently surpassing both time series and multi-modal baselines by up to 8.9% of AUC (compared to TimeCAP on Weather dataset). This underscores the advantage of TimeXL, which synergizes multi-modal time series encoder with language agents to enhance interpretability and thus predictive performance in multi-modal time series.
@@ 4.3 Explainable Multi-modal Prototypes
Next, we present the explainable multi-modal prototypes rendered by TimeXL, which establishes the case-based reasoning process. Figure 3 shows a subset of time series and text prototypes learned on the weather dataset. The time series prototypes demonstrate the typical temporal patterns aligned with different real-world weather conditions (i.e., rain and not rain). For example, a constant or decreasing humidity at a moderate level, combined with high and steady air pressure, typically indicates a non-rainy scenario. The consistent wind direction is also a sign of mild weather conditions. On the contrary, high humidity, low and fluctuating pressure, along with variable winds typically reveal an unstable weather system ahead. In addition to time series, the text prototypes also highlight consistent semantic patterns for different weather conditions, such as the channel-specific (e.g., drier air moving into the area, strengthening of high-pressure system) and overall (e.g., a likelihood of dry weather) descriptions of weather activities. In Appendix C.1, we also present more multi-modal prototypes for the weather dataset in Figure 8, for other datasets in Figures 9 and 10. The results validate that TimeXL provides coherent and informative prototypes from the exploitation of time series and its real-world contexts, which facilitates both prediction and explanation.

### v2

@@ 3.3.2 Iterative Context Refinement via Reflective Feedback
Given the prediction $\boldsymbol{\hat{y}}_{\mathrm{LLM}}$ generated by the prediction agent $\mathcal{M}_{\mathrm{pred}}$, the reflection agent $\mathcal{M}_{\mathrm{refl}}$ aims to understand the reasoning behind the implicit prediction logic of $\mathcal{M}_{\mathrm{pred}}$. Specifically, it generates a reflective feedback, $\mathrm{Refl}$, by analyzing the input text $\boldsymbol{s}$ and its prediction $\boldsymbol{\hat{y}}_{\mathrm{LLM}}$, against the ground truth $\boldsymbol{y}$, to provide actionable insights for refinement, i.e., $\mathrm{Refl}=\mathcal{M}_{\mathrm{refl}}(\boldsymbol{y},\boldsymbol{\hat{y}}_{\mathrm{LLM}},\boldsymbol{s})$. Guided by the feedback, the refinement agent $\mathcal{M}_{\mathrm{refine}}$ refines the previous text $\boldsymbol{s}_{i}$ into $\boldsymbol{s}_{i+1}$ by selecting and emphasizing the most relevant content, ensuring that important patterns are appropriately contextualized, which is similar to how a domain expert would perform, i.e., $\boldsymbol{s_{i+1}}=\mathcal{M}_{\mathrm{refine}}(\mathrm{Refl},\boldsymbol{s}_{i})$. The prompts for querying $\mathcal{M}_{\mathrm{refl}}$ and $\mathcal{M}_{\mathrm{refine}}$ are provided in Figures 1617 1819, and discussed in Appendix D.
Algorithm 1 Iterative Optimization Loop of TimeXL
Inputs: Multi-modal time series $(\boldsymbol{x},\boldsymbol{s},\boldsymbol{y})$ with $\mathcal{D}_{\mathrm{train}}$, $\mathcal{D}_{\mathrm{val}}$, $\mathcal{D}_{\mathrm{test}}$, prototype-based encoder $\mathcal{M}_{\mathrm{enc}}$, prediction agent $\mathcal{M}_{\mathrm{pred}}$, reflection agent $\mathcal{M}_{\mathrm{refl}}$, refinement agent $\mathcal{M}_{\mathrm{refine}}$, fusion parameter $\alpha$, max iteration $\tau$, improvement check $\mathrm{Eval(\cdot)}$ Training & Validation:
Initialize $\boldsymbol{s}_{0}=\boldsymbol{s}$, $i=0$, $\mathcal{M}_{\mathrm{enc}}^{\ast}\leftarrow\varnothing$, $\mathrm{Refl}^{\ast}\leftarrow\varnothing$
while $i<\tau$ do
       Train $\mathcal{M}_{\mathrm{enc}}$ using $\mathcal{D}_{\mathrm{train,i}}=\{(\boldsymbol{x},\boldsymbol{s}_{i},\boldsymbol{y}),\cdots\}$
$\boldsymbol{\hat{y}}_{\mathrm{enc}},\mathbf{expl}_{\boldsymbol{s}_{i}}=\mathcal{M}_{\mathrm{enc}}(\boldsymbol{x},\boldsymbol{s}_{i})$
$\boldsymbol{\hat{y}}_{\mathrm{LLM}}=\mathcal{M}_{\mathrm{pred}}(\boldsymbol{s}_{i},\mathbf{expl}_{\boldsymbol{s}_{i}})$
$\boldsymbol{\hat{y}}=\alpha\boldsymbol{\hat{y}}_{\mathrm{enc}}+(1-\alpha)\boldsymbol{\hat{y}}_{\mathrm{LLM}}$
$\mathrm{Refl}=\mathcal{M}_{\mathrm{refl}}(\boldsymbol{y},\boldsymbol{\hat{y}}_{\mathrm{LLM}},\boldsymbol{s}_{i})$
$\boldsymbol{s_{i+1}}=\mathcal{M}_{\mathrm{refine}}(\mathrm{Refl},\boldsymbol{s}_{i})$
if $\mathrm{Eval}(\mathrm{relf},\mathcal{D}_{\mathrm{val}})$ pass then
            $\mathcal{M}_{\mathrm{enc}}^{\ast}\leftarrow\mathcal{M}_{\mathrm{enc}}$ , $\mathrm{Refl}^{\ast}\leftarrow\mathrm{Refl}$
@@
return $\mathcal{M}_{\mathrm{enc}}^{\ast}$, $\mathrm{Refl}^{\ast}$ Testing:
$\boldsymbol{s^{\prime}}=\mathcal{M}_{\mathrm{refine}}(\mathrm{Refl}^{\ast},\boldsymbol{s})$ for $\mathcal{D}_{\mathrm{test}}$
$\boldsymbol{\hat{y}}_{\mathrm{enc}},\mathbf{expl}_{\boldsymbol{s^{\prime}}}=\mathcal{M}^{\ast}_{\mathrm{enc}}(\boldsymbol{x},\boldsymbol{s^{\prime}})$
$\boldsymbol{\hat{y}}_{\mathrm{LLM}}=\mathcal{M}_{\mathrm{pred}}(\boldsymbol{s^{\prime}},\mathbf{expl}_{\boldsymbol{s^{\prime}}})$
$\boldsymbol{\hat{y}}=\alpha\boldsymbol{\hat{y}}_{\mathrm{enc}}+(1-\alpha)\boldsymbol{\hat{y}}_{\mathrm{LLM}}$
We finally integrate the refinement via reflection into the optimization loop of our proposed TimeXL, which is summarized in Algorithm 1. Once the textual context is improved, it is used to retrain the multi-modal prototype-based encoder $\mathcal{M}_{\mathrm{enc}}$ for the next iteration. As such, the explanation (e.g., quality of the prototypes) and predictive performance of $\mathcal{M}_{\mathrm{enc}}$ can be improved through this iterative process. Consequently, the prediction agent $\mathcal{M}_{\mathrm{pred}}$ could yield better prediction with more informative inputs, further enhancing the accuracy of $\boldsymbol{\hat{y}}$. We evaluate the predictive performance of LLM and save the encoder and reflection when an improvement is observed ($\mathrm{Eval(\cdot)}$ pass) when applied to the validation set. In the testing phase, we use the reflection $\mathrm{Refl}$ generated in the iteration with the best validation performance to guide $\mathcal{M}_{\mathrm{refine}}$ for context refinement, mimicking how an optimized deep model is applied to testing data.
@@ 4 Experiments
Table 1: The F1 score (F1) and AUROC (AUC) for TimeXL and state-of-the-art baselines on multi-modal time series datasets.
@@
DLinear [93] 0.540 0.660 0.255 0.485 0.393 0.500 0.419 0.388
Autoformer [94] 0.546 0.590 0.565 0.747 0.774 0.918 0.683 0.825
Crossformer [95] 0.500 0.594 0.571 0.775 0.924 0.984 0.737 0.913
TimesNet [96] 0.494 0.594 0.538 0.756 0.794 0.867 0.765 0.944
iTransformer [97] 0.541 0.650 0.600 0.783 0.861 0.931 0.791 0.963
TSMixer [98] 0.488 0.534 0.465 0.689 0.770 0.797 0.808 0.931
TimeMixer [99] 0.577 0.658 0.571 0.776 0.822 0.887 0.824 0.935
FreTS [100] 0.623 0.688 0.546 0.737 0.887 0.950 0.751 0.762
PatchTST [5] 0.592 0.675 0.604 0.795 0.841 0.934 0.695 0.928
LLMTime [101] 0.587 0.657 0.519 0.643 0.802 0.817 0.769 0.803
PromptCast [79] 0.499 0.365 0.418 0.607 0.727 0.768 0.696 0.871
OFA [66] 0.501 0.606 0.512 0.745 0.774 0.879 0.851 0.977
FSCA [73] 0.563 0.647 0.592 0.790 0.820 0.891 0.872 0.977
Time-LLM [69] 0.613 0.699 0.589 0.792 0.671 0.864 0.733 0.912
TimeCMA [72] 0.636 0.731 0.559 0.727 0.729 0.828 0.693 0.843
MM-iTransformer [47] 0.608 0.689 0.605 0.793 0.926 0.986 0.901 0.990
MM-PatchTST [47] 0.621 0.718 0.619 0.812 0.863 0.968 0.780 0.929
TimeCAP [81] 0.668 0.742 0.611 0.801 0.954 0.983 0.942 0.988
@@ 4.1 Experimental Setup
Datasets. We evaluate methods on four multi-modal time series datasets from three different real-world domains, including weather, finance, and healthcare. The detailed data statistics are summarized in Table 3 of Appendix A.1. The weather dataset contains meteorological reports and the hourly time series records of temperature, humidity, air pressure, wind speed, and wind direction in New York City. The task is to predict if it will rain in the next 24 hours, given the last 24 hours of weather records and summary. The finance dataset contains the daily record of the raw material prices together with 14 related indices from January 2017 to July 2024. Given the last 5 business days of stock price data and news, the task is to predict if the target price will exhibit an increasing, decreasing, or neutral trend on the next business day. The healthcare datasets include Test-Positive (TP) and Mortality (MT), both comprising weekly records related to influenza activity. TP contains the number of respiratory specimens testing positive for Influenza A and B, while MT reports deaths caused by influenza and pneumonia. For both datasets, the task is to predict whether the respective target ratio, either the test-positive rate or the mortality ratio in the upcoming week will exceed the historical average, based on records and summaries from the previous 20 weeks.
Baselines, Evaluation Metrics, and Setup We compare TimeXL with state-of-the-art baselines for time series prediction, including Autoformer [94], Dlinear [93], Crossformer [95], TimesNet [96], PatchTST [5], iTransformer [97], FreTS [100], TSMixer [98], TimeMixer [99] (classification implemented by TSlib [96]) and LLM-based methods like LLMTime [101], PromptCast [79], OFA [66], Time-LLM [69], TimeCMA [72] and FSCA [73], where LLMTime and PromptCast don’t need fine-tuning. While some methods are primarily used for time series prediction with continuous values, they can be easily adapted for discrete value prediction. We also evaluate the multi-modal time series methods. Besides the Time-LLM, TimeCMA and FSCA where input text is used for embedding reprogramming and alignment, we also evaluate Multi-modal PatchTST and Multi-modal iTransformer from [47], as well as TimeCAP [81]. We evaluate the discrete prediction via F1 score and AUROC (AUC) score, due to label imbalance in real-world time series datasets. All datasets are split 6:2:2 for training, validation, and testing. All results are averaged over five experimental runs. We alternate different embedding methods for texts based on its average length (Sentence-BERT [92] for finance dataset and BERT [91] for the others). More details are provided in Appendix A.2A.3A.4.
@@ 4.2 Performance Evaluation
The results of predictive performance are shown in Table 1. It is notable that multi-modal methods generally outperform time series methods across all datasets. These methods include LLM methods (e.g., Time-LLM [69], TimeCMA [72], FSCA [73]) that leverage text embeddings to enhance time series predictions. Moreover, the multi-modal variants (MM-iTransformer and MM-PatchTST [47]) improve the performance of state-of-the-art time series methods, suggesting the benefits of integrating real-world contextual information. Besides, TimeCAP [81] integrates the predictions from both modalities, further improving the predictive performance. TimeXL constantly achieves the highest F1 and AUC scores, consistently surpassing both time series and multi-modal baselines by up to 8.9% of AUC (compared to TimeCAP on Weather dataset). This underscores the advantage of TimeXL, which synergizes multi-modal time series encoder with language agents to enhance interpretability and thus predictive performance in multi-modal time series.
@@ 4.3 Explainable Multi-modal Prototypes
Next, we present the explainable multi-modal prototypes rendered by TimeXL, which establishes the case-based reasoning process. Figure 3 shows a subset of time series and text prototypes learned on the weather dataset. The time series prototypes demonstrate the typical temporal patterns aligned with different real-world weather conditions (i.e., rain and not rain). For example, a constant or decreasing humidity at a moderate level, combined with high and steady air pressure, typically indicates a non-rainy scenario. The consistent wind direction is also a sign of mild weather conditions. On the contrary, high humidity, low and fluctuating pressure, along with variable winds typically reveal an unstable weather system ahead. In addition to time series, the text prototypes also highlight consistent semantic patterns for different weather conditions, such as the channel-specific (e.g., drier air moving into the area, strengthening of high-pressure system) and overall (e.g., a likelihood of dry weather) descriptions of weather activities. In Appendix C.1, we also present more multi-modal prototypes for the weather dataset in Figure 9, for the finance dataset in Figure 10, and for healthcare datasets in Figures 11 and 12. TimeXL provides coherent and informative prototypes from the exploitation of time series and its real-world contexts, which facilitates both prediction and explanation.